OpenStudy (anonymous):

The ratio of men to women in a certain factory is 3 to 4. There are 204 men. How many workers are there?

OpenStudy (calculusxy):

Answer choices?

OpenStudy (anonymous):

It's 3 men to 4 women. 3 out of 7 are men. 204 = 3/7 of the total 68 = 1/7 of the total The total is 476.

OpenStudy (anonymous):

3/4=204/x 3x=816 x=272 (that is the number of female employees) 204+272=476 total number

OpenStudy (anonymous):

Did this help?

OpenStudy (calculusxy):

I got 676 somehow.

OpenStudy (anonymous):

lol, no. :)

OpenStudy (anonymous):

it helped me bro thanks (:

OpenStudy (anonymous):

No problem! :D

OpenStudy (calculusxy):

Proportions. 3/4=204/x , where x is women x=272 Add and get 576 (not 676)

OpenStudy (anonymous):

? You are 100 people off

OpenStudy (calculusxy):

Yes but I want to know how.

OpenStudy (anonymous):

476 was correct

OpenStudy (anonymous):

Did you see me break it down hon? YAY JCP!! :D

OpenStudy (calculusxy):

Yes I added wrong . My mistake

OpenStudy (anonymous):

hehe - don't stress man

OpenStudy (calculusxy):

No. I'm fine :)

OpenStudy (anonymous):

I used to do that all the time

OpenStudy (anonymous):

A medal for your worries

OpenStudy (anonymous):

@Saphira_Shapiro The perimeter of a rectangle is twice the sum of its length and its width. The perimeter is 28 meters and its length is 2 meters more than twice its width.

OpenStudy (anonymous):

Let's see.........

OpenStudy (anonymous):

thanks bro

OpenStudy (mathmale):

I can help, but I'm tied up at the moment helping 2 other peoople. It's up to you wehther to wait for me or to move on.

OpenStudy (anonymous):

i'll wait

OpenStudy (anonymous):

No prob Jc

OpenStudy (anonymous):

use L for length and W for width: Perimeter is: 2*L + 2*W = 28. Length is 2 meters more than twice the width: L = 2*W+2. Use the above for length in the perimeter: 2*(2*W+2) + 2*W = 28.

OpenStudy (anonymous):

The perimeter is (=) twice (two times) the sum of its length and width (L+W). P=2(L+W) The perimeter is (=) 28 meters and its length is 2 meters more (+2) than twice (times 2) its width. What is the length? P=28 L=2+2W So you know that P=28. Plug that into the first equation: P=2(L+W) P=28 28=2(L+W) then simplify by distributing 28=2L+2W You also know that L=2+2W, so substitute that for L

OpenStudy (anonymous):

It cut me off so I just pasted the rest

OpenStudy (anonymous):

I'm a chic - lol

OpenStudy (mathmale):

JC and Chic: Hello. Do you two have this problem solution under control yourselves, or do y ou need pointers?

OpenStudy (anonymous):

so the answer is 28=2L+2W /